Shift report 20000824 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 24 2000 - 15:57:02 EDT

  • Next message: http account: "Shift report 20000825 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Aug 24 15:57:02 2000, Shift: 08:00-16:00
    Supervisor: HI, with: BA
    Log book#5 pages 8 - 8
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam until 12:56.  Short opening.
    2.  continued run2423.  and run2424
      We have about 400k trigger 1 in run2423 with 2424.
      The ratio of trigger 1 to collision is about 1.3/1.
    3.  T1 is not working.  T1's anode voltage does not hold.
    By Ramiro's request, the field voltage of T1 is raised to
    700V.  T1 anode voltage does not trip, but there are very 
    tracks in T1.
      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 24 2000 - 16:01:27 EDT